# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ B.4 Effect of OTGP flow
We show (B.15) in this section.
@@
Under the same assumption as Theorem 4.4,
@@
We observe from (B.106) that the derivative in (B.105) can be decomposed into three terms
$\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}=(\text{I})+(\text{II})+(\text{III})$ (B.107)
@@
Step 1. By Lemma A.12, $|\nabla_{x}V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}(s,y)|\leq C(1+|x|)$. We temporally fix $s$, $x$, $\alpha=v\alpha^{\tau}(s,x)+(1-v)\alpha^{\mu^{\tau},*}(s,x)$, and $p=-\nabla_{x}V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}(s,x)$. Taking derivative of $\nabla_{\alpha}^{2}H(s,x,\mu^{\tau}_{s},\alpha,p)$ with respect to $\tau$, we have
@@
$\leq\lim_{\Delta\tau\to 0}\dfrac{1}{|\Delta\tau|}C\,W_{2}(\mu^{\tau+\Delta\tau}_{s},\mu^{\tau}_{s})\,(1+|x|)=C(1+|x|)\,\beta_{\mu}\,W_{2}(\rho^{\mu^{\tau},\alpha^{\tau}}_{s},\mu^{\tau}_{s})\leq C(1+|x|)\,\beta_{\mu}.$
Therefore, term $(\text{I})$ satisfies
@@
where $\alpha$ and $p$ are evaluated at $v\alpha^{\tau}(s,x)+(1-v)\alpha^{\mu^{\tau},*}(s,x)$ and $\nabla_{x}V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}(s,x)$ in the first line. The second inequality is due to (B.108), and the third inequality is because $\alpha^{\tau}-\alpha^{\mu^{\tau},*}\in\mathcal{C}$.
Step 2. For $(\text{II})$, we want to show
$\Big|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\partial_{x_{1}}V^{\mu^{\tau},\alpha}(s,x)\Big|\leq C\beta_{\mu}(1+|x|)$
first (cf. (A.20)). Let $\Delta\tau,\delta\in\mathbb{R}$, denote $\mu=\mu^{\tau}$, $\alpha=\alpha^{\tau}$ and $\mu^{\prime}=\mu^{\tau+\Delta\tau}$ for simplicity. Let $x\in\mathbb{R}^{d}$ and $x^{\delta}:=x+\delta e_{1}$. Then
@@
We define $b_{\alpha}(t,x,\mu_{t}):=b(t,x,\mu_{t},\alpha(t,x))$ and $f_{\alpha}$ similarly. Let $x_{t},x^{\delta}_{t},x^{\prime}_{t},x^{\delta\prime}$ be the state processes that start at $x,x^{\delta},x,x^{\delta}$, with drifts
$b_{t}:=b_{\alpha}(t,x_{t},\mu_{t}),~b^{\delta}_{t}:=b_{\alpha}(t,x^{\delta}_{t},\mu_{t}),~b^{\prime}_{t}:=b_{\alpha}(t,x_{t}^{\prime},\mu_{t}^{\prime}),~b^{\delta\prime}_{t}:=b_{\alpha}(t,x^{\delta\prime}_{t},\mu_{t})$
and diffusions $\sigma_{t},\sigma^{\delta}_{t},\sigma^{\prime}_{t},\sigma^{\delta\prime}_{t}$ defined similarly. We also define $f_{t},f^{\delta}_{t},f^{\prime}_{t},f^{\delta\prime}_{t}$ in a similar manner. We also assume that the four state processes $x_{t},x^{\delta}_{t},x^{\prime}_{t},x^{\delta\prime}$ are driven by the same Brownian motion. Then, by definition (2.4),
@@
Note that
$f^{\delta}_{t}-f_{t}$ $=f_{\alpha}(t,x^{\delta}_{t},\mu_{t})-f_{\alpha}(t,x_{t},\mu_{t})=\int_{0}^{1}(x^{\delta}_{t}-x_{t})^{\top}\,\nabla_{x}f_{\alpha}(t,(1-u)x_{t}+ux^{\delta}_{t},\mu_{t})\,\mathrm{d}u$ (B.112)
@@
We subtract the two equations in (B.112), and note that $\nabla_{x}f_{\alpha}=\nabla_{x}f+\nabla_{x}\alpha^{\top}\nabla_{\alpha}f$ is Lipschitz in $x$ and $\mu_{t}$ and grows at most linearly in $|x|$, we get
@@
Taking expectation, we get
@@
$\leq C\Big[(1+|x|)\,(|x-x^{\delta}|\,\mathcal{W}_{2}(\mu,\mu^{\prime}))+|x-x^{\delta}|\,\mathcal{W}_{2}(\mu,\mu^{\prime})\Big]\leq C\,\delta\,(1+|x|)\,\mathcal{W}_{2}(\mu,\mu^{\prime}),$
where we used Hölder’s inequality in the first inequality, and Grönwall’s inequalities (A.1), (A.2), (A.9) from Section A.1 in the second inequality. Similarly, we can show
$\mathbb{E}\Big[\left|(g(x^{\delta\prime}_{T},\mu^{\prime}_{T})-g(x^{\prime}_{T},\mu^{\prime}_{T}))+(g(x^{\delta}_{T},\mu_{T})-g(x_{T},\mu_{T}))\right|\Big]\leq C\,\delta\,(1+|x|)\,\mathcal{W}_{2}(\mu,\mu^{\prime}).$ (B.115)
Substituting (B.114) and (B.115) into (B.111), we get
$\Big|\big(V^{\mu^{\prime},\alpha}(s,x^{\delta})-V^{\mu^{\prime},\alpha}(s,x)\big)-\big(V^{\mu,\alpha}(s,x^{\delta})-V^{\mu,\alpha}(s,x)\big)\Big|\leq C\,\delta\,(1+|x|)\,\mathcal{W}_{2}(\mu,\mu^{\prime}).$ (B.116)
Substituting (B.116) into (B.110), we get
$\bigg|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\partial_{x_{1}}V^{\mu^{\tau},\alpha}(s,x)\bigg|\leq C(1+|x|)\lim_{\Delta\tau\to 0}\frac{1}{|\Delta\tau|}\mathcal{W}_{2}(\mu^{\tau},\mu^{\tau+\Delta\tau})=C\beta_{\mu}(1+|x|)\,\mathcal{W}_{2}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}).$ (B.117)
where we used Lemma A.15 in the equality. Applying (B.117) in each dimension, we get
$\bigg|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\nabla_{x}V^{\mu^{\tau},\alpha}(s,x)\bigg|\leq C\beta_{\mu}(1+|x|)\,\mathcal{W}_{2}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})\leq C\beta_{\mu}(1+|x|).$
Therefore,
@@
Step 3. We estimate $(\text{III})$ in this step. We define
$q^{\tau}(t,x)=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\log\rho^{\mu^{\tau},\alpha}(t,x)\Big|_{\alpha=\alpha^{\tau}}.$
We will show $\left|q^{\tau}(t,x)\right|\leq C\beta_{\mu}(1+|x|^{2})$. We will use the short notation $b$ and $D$, and $\rho$ to denote $b(t,x,\mu^{\tau}_{t},\alpha^{\tau}(t,x))$, $D(t,x,\mu_{t})$ and $\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)$ for simplicity. Since $\rho$ satisfies the FP equation (2.6),
$\quad\partial_{t}\log\rho=\partial_{t}\rho/\rho=-\nabla_{x}\cdot(b\rho)/\rho+\nabla_{x}^{2}:(D\rho)/\rho$ (B.119)
$=-\nabla_{x}\cdot b+\nabla_{x}^{2}:D-b^{\top}\nabla_{x}\log\rho+2(\nabla_{x}\cdot D)^{\top}\nabla_{x}\log\rho+\operatorname{Tr}[D(\nabla_{x}^{2}\log\rho+\nabla_{x}\log\rho\,\nabla_{x}\log\rho^{\top})],$
where we use $\nabla_{x}^{2}:$ to denote the trace inner product between the Hessian operator and a matrix valued function. Taking derivative in $\tau$, we get
$\partial_{t}q^{\tau}$ $=-\nabla_{x}\cdot\partial_{\tau}b+\nabla_{x}^{2}:\partial_{\tau}D-\partial_{\tau}b^{\top}\nabla_{x}\log\rho+2(\nabla_{x}\cdot\partial_{\tau}D)^{\top}\nabla_{x}\log\rho-b^{\top}\nabla_{x}q^{\tau}+2(\nabla_{x}\cdot D)\nabla_{x}q^{\tau}$ (B.120)
@@
$=:a_{q}+b_{q}^{\top}\nabla_{x}q^{\tau}+\operatorname{Tr}[D\nabla_{x}^{2}q^{\tau}]$
(B.120) is a linear parabolic equation for $q^{\tau}$ with initial condition $q^{\tau}(0,x)=0$. Using the assumptions, we have $|a_{q}|\leq C(1+|x|^{2})$, $b_{q}\leq C(1+|x|)$. By the Lipschitz condition of $b,\nabla_{x}b,\nabla_{x}D,\nabla_{x}^{2}D$ in $\mu_{t}$ and Lemma A.15, we have
@@
Therefore, applying standard maximum principle with a quadratic barrier function [36] to the PDE $\partial_{t}q^{\tau}=a_{q}+b_{q}^{\top}\nabla_{x}q^{\tau}+\operatorname{Tr}[D\nabla_{x}^{2}q^{\tau}]$ with initial condition $q^{\tau}(0,x)=0$, we get $|q^{\tau}(t,x)|\leq C\beta_{\mu}(1+|x|^{2})$, which implies $\tfrac{\mathrm{d}}{\mathrm{d}\tau}\rho^{\mu^{\tau},\alpha}(s,x)\big|_{\alpha=\alpha^{\tau}}\leq C\beta_{\mu}\rho^{\mu^{\tau},\alpha}(s,x)$. Therefore,
$(\text{III})$ $\leq C\beta_{\mu}\int_{0}^{T}\int_{\mathbb{R}^{d}}\int_{0}^{1}\int_{0}^{u}(1+|x|)\left|\alpha^{\tau}(s,x)-\alpha^{\mu^{\tau},*}(s,x)\right|^{2}\,\mathrm{d}v\,\mathrm{d}u\,\Big|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\rho^{\mu^{\tau},\alpha}(s,x)\big|_{\alpha=\alpha^{\tau}}\Big|\,\mathrm{d}x\,\mathrm{d}s$ (B.121)
@@
Finally, combining (B.109), (B.118), and (B.121), we get
$\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}\leq C\beta_{\mu}\left\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\right\|_{\mu^{\tau},\alpha^{\tau}}^{2}.$

### v2

@@ B.4 Effect of OTGP flow
In this section, we show (B.15), which is stated as Lemma B.3 below.
@@
Under the conditions of Theorem 4.4,
@@
$\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}=(\text{I})+(\text{II})+(\text{III}).$ (B.102)
$(\text{I})$ addresses the $\tau$-dependence through the third argument of $\nabla_{\alpha}^{2}H$. $(\text{II})$ addresses the $\tau$-dependence through the first superscript of $\nabla_{x}V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}$ in the fifth argument of $\nabla_{\alpha}^{2}H$. $(\text{III})$ addresses the $\tau$-dependence through the first superscript of the density $\rho^{\mu^{\tau},\alpha^{\tau}}$. Note that we use the notation $V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}$ instead of $V^{\mu^{\tau},*}$ to clearly distinguish the $\tau$-dependence of the distribution and the control components.
In the following context, we estimate each of the three parts separately.
Step 1. By Lemma A.12, $|\nabla_{x}V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}(s,x)|\leq C(1+|x|)$. For notational simplicity, we temporarily fix $s$, $x$, $\alpha:=v\alpha^{\tau}(s,x)+(1-v)\alpha^{\mu^{\tau},*}(s,x)$ and $p:=-\nabla_{x}V^{\mu^{\tau},\alpha^{\mu^{\tau},*}}(s,x)$. Differentiating $\nabla_{\alpha}^{2}H(s,x,\mu^{\tau}_{s},\alpha,p)$ with respect to $\tau$ yields
@@
$\leq\lim_{\Delta\tau\to 0}\dfrac{1}{|\Delta\tau|}C\,W_{2}(\mu^{\tau+\Delta\tau}_{s},\mu^{\tau}_{s})\,(1+|x|)=C(1+|x|)\,\beta_{\mu}\,W_{2}(\rho^{\mu^{\tau},\alpha^{\tau}}_{s},\mu^{\tau}_{s})\leq C(1+|x|)\,\beta_{\mu},$
where the last equality follows from Lemma A.15. Here, we are also using the uniform (in $\tau$ and $t$) boundedness of $W_{2}(\rho^{\mu^{\tau},\alpha^{\tau}}_{t},\mu^{\tau}_{t})\leq W_{2}(\rho^{\mu^{\tau},\alpha^{\tau}}_{t},\delta_{0})+W_{2}(\delta_{0},\mu^{\tau}_{t})$, which is implied by $\mu^{\tau}\in\mathcal{M}$ and the Aronson-type bound (4.1). Therefore, term $(\text{I})$ satisfies
@@
where the last inequality is due to $\alpha^{\tau}-\alpha^{\mu^{\tau},*}\in\mathcal{C}$.
Step 2. Motivated by (A.20), we first show
$\Big|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\partial_{x_{1}}V^{\mu^{\tau},\alpha}(s,x)\Big|\leq C\beta_{\mu}(1+|x|).$
Let $\Delta\tau,\delta\in\mathbb{R}$, denote $\mu:=\mu^{\tau}$, $\alpha:=\alpha^{\tau}$, $\mu^{\prime}:=\mu^{\tau+\Delta\tau}$ and $x^{\delta}:=x+\delta e_{1}$ for $x\in\mathbb{R}^{d}$.
@@
Denote $b_{\alpha}(t,x,\mu_{t}):=b(t,x,\mu_{t},\alpha(t,x))$ and let $f_{\alpha}$ be similarly defined. Define $x_{t},x^{\delta}_{t},x^{\prime}_{t},x^{\delta\prime}_{t}$ as state processes driven by the same Brownian motion that have initial conditions $x,x^{\delta},x,x^{\delta}$ at time $s$, with respective drifts
$b_{t}:=b_{\alpha}(t,x_{t},\mu_{t}),~b^{\delta}_{t}:=b_{\alpha}(t,x^{\delta}_{t},\mu_{t}),~b^{\prime}_{t}:=b_{\alpha}(t,x_{t}^{\prime},\mu_{t}^{\prime}),~b^{\delta\prime}_{t}:=b_{\alpha}(t,x^{\delta\prime}_{t},\mu_{t}),$
and diffusions $\sigma_{t},\sigma^{\delta}_{t},\sigma^{\prime}_{t},\sigma^{\delta\prime}_{t}$ that are similarly defined. The processes $f_{t},f^{\delta}_{t},f^{\prime}_{t},f^{\delta\prime}_{t}$ are defined in a similar manner. By definition (2.4),
@@
By the mean value theorem,
$f^{\delta}_{t}-f_{t}$ $=f_{\alpha}(t,x^{\delta}_{t},\mu_{t})-f_{\alpha}(t,x_{t},\mu_{t})=\int_{0}^{1}(x^{\delta}_{t}-x_{t})^{\top}\,\nabla_{x}f_{\alpha}(t,(1-u)x_{t}+ux^{\delta}_{t},\mu_{t})\,\mathrm{d}u,$ (B.107)
@@
By Assumption 4.1, $\nabla_{x}f_{\alpha}=\nabla_{x}f+\nabla_{x}\alpha^{\top}\nabla_{\alpha}f$ is Lipschitz in $x,\mu$ and grows at most linearly in $|x|$. Subtracting the two equations in (B.107) yields
@@
Taking expectations on both sides yields
@@
$\leq C\Big[(1+|x|)\,|x-x^{\delta}|\,\mathcal{W}_{2}(\mu,\mu^{\prime})+|x-x^{\delta}|\,\mathcal{W}_{2}(\mu_{t},\mu^{\prime}_{t})\Big]$
$\leq C\,\delta\,[(1+|x|)\,\mathcal{W}_{2}(\mu,\mu^{\prime})+W_{2}(\mu^{\prime}_{t},\mu_{t})],$
where Grönwall’s inequalities (A.1), (A.2), (A.9) are applied. Similarly,
$\mathbb{E}\Big[\left|(g(x^{\delta\prime}_{T},\mu^{\prime}_{T})-g(x^{\prime}_{T},\mu^{\prime}_{T}))+(g(x^{\delta}_{T},\mu_{T})-g(x_{T},\mu_{T}))\right|\Big]\leq C\,\delta\,[(1+|x|)\,\mathcal{W}_{2}(\mu,\mu^{\prime})+W_{2}(\mu^{\prime}_{T},\mu_{T})].$ (B.110)
Substituting (B.109) and (B.110) into (B.106) yields
$\Big|\big(V^{\mu^{\prime},\alpha}(s,x^{\delta})-V^{\mu^{\prime},\alpha}(s,x)\big)-\big(V^{\mu,\alpha}(s,x^{\delta})-V^{\mu,\alpha}(s,x)\big)\Big|\leq C\,\delta\,[(1+|x|)\,\mathcal{W}_{2}(\mu,\mu^{\prime})+W_{2}(\mu^{\prime}_{T},\mu_{T})].$ (B.111)
Substituting (B.111) into (B.105) yields
$\quad\bigg|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\partial_{x_{1}}V^{\mu^{\tau},\alpha}(s,x)\bigg|\leq C\Big[(1+|x|)\lim_{\Delta\tau\to 0}\frac{1}{|\Delta\tau|}\mathcal{W}_{2}(\mu^{\tau},\mu^{\tau+\Delta\tau})+\lim_{\Delta\tau\to 0}\frac{1}{|\Delta\tau|}W_{2}(\mu^{\tau}_{T},\mu^{\tau+\Delta\tau}_{T})\Big]$ (B.112)
$=C\beta_{\mu}[(1+|x|)\,\mathcal{W}_{2}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})+\beta_{\mu}W_{2}(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T})]\leq C\beta_{\mu}(1+|x|)$
where Lemma A.15 is applied. Repeating the argument (B.112) for each dimension yields $\bigg|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\nabla_{x}V^{\mu^{\tau},\alpha}(s,x)\bigg|\leq C\beta_{\mu}(1+|x|)$. By Assumption 4.1, $\alpha^{\tau}-\alpha^{\mu^{\tau},*}\in\mathcal{C}$ and previous estimations,
@@
Step 3. We estimate $(\text{III})$. Define
$q^{\tau}(t,x):=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\log\rho^{\mu^{\tau},\alpha}(t,x)\Big|_{\alpha=\alpha^{\tau}}.$
We claim that: $\left|q^{\tau}(t,x)\right|\leq C\beta_{\mu}(1+|x|^{2})$.
We use shorthand notations $b$, $D$, $\rho$ to denote $b(t,x,\mu^{\tau}_{t},\alpha^{\tau}(t,x))$, $D(t,x,\mu^{\tau}_{t})$, $\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)$. Since $\rho$ satisfies the FP equation (2.6),
$\quad\partial_{t}\log\rho=\partial_{t}\rho/\rho=-\frac{\nabla_{x}\cdot(b\rho)}{\rho}+\frac{\nabla_{x}^{2}:(D\rho)}{\rho}$ (B.114)
$=-\nabla_{x}\cdot b+\nabla_{x}^{2}:D-b^{\top}\nabla_{x}\log\rho+2(D\cdot\nabla_{x})^{\top}\nabla_{x}\log\rho+\operatorname{Tr}[D(\nabla_{x}^{2}\log\rho+\nabla_{x}\log\rho\,\nabla_{x}\log\rho^{\top})],$
where $\nabla_{x}^{2}:$ denotes the matrix inner product (i.e., $\langle A,B\rangle:=\operatorname{Tr}(A^{\top}B)$) between the Hessian operator and a matrix-valued function. Differentiating with respect to $\tau$ yields
$\partial_{t}q^{\tau}$ $=-\nabla_{x}\cdot\partial_{\tau}b+\nabla_{x}^{2}:\partial_{\tau}D-\partial_{\tau}b^{\top}\nabla_{x}\log\rho+2(\partial_{\tau}D\cdot\nabla_{x})^{\top}\nabla_{x}\log\rho-b^{\top}\nabla_{x}q^{\tau}+2(D\cdot\nabla_{x})\nabla_{x}q^{\tau}$ (B.115)
@@
$=:a_{q}+b_{q}^{\top}\nabla_{x}q^{\tau}+\operatorname{Tr}[D\nabla_{x}^{2}q^{\tau}],$
which is a linear parabolic equation for $q^{\tau}$ with initial condition $q^{\tau}(0,x)=0$. By the Lipschitz condition of $b,\nabla_{x}b,\nabla_{x}D,\nabla_{x}^{2}D$ in $\mu$ and Lemma A.15, we have
@@
By the logarithmic Aronson bounds, $|a_{q}|\leq C\beta_{\mu}(1+|x|^{2})$, $b_{q}\leq C\beta_{\mu}(1+|x|)$. Applying standard maximum principle with a quadratic barrier function [36] to the PDE $\partial_{t}q^{\tau}=a_{q}+b_{q}^{\top}\nabla_{x}q^{\tau}+\operatorname{Tr}[D\nabla_{x}^{2}q^{\tau}]$ with initial condition $q^{\tau}(0,x)=0$ yields $|q^{\tau}(t,x)|\leq C\beta_{\mu}(1+|x|^{2})$, which implies $\tfrac{\mathrm{d}}{\mathrm{d}\tau}\rho^{\mu^{\tau},\alpha}(s,x)\big|_{\alpha=\alpha^{\tau}}\leq C\beta_{\mu}(1+|x|^{2})\rho^{\mu^{\tau},\alpha^{\tau}}(s,x)$. By (B.94),
$(\text{III})$ $\leq\int_{0}^{T}\int_{\mathbb{R}^{d}}\int_{0}^{1}\int_{0}^{u}(1+|x|)\left|\alpha^{\tau}(s,x)-\alpha^{\mu^{\tau},*}(s,x)\right|^{2}\,\mathrm{d}v\,\mathrm{d}u\,\Big|\dfrac{\mathrm{d}}{\mathrm{d}\tau}\rho^{\mu^{\tau},\alpha}(s,x)\big|_{\alpha=\alpha^{\tau}}\Big|\,\mathrm{d}x\,\mathrm{d}s$ (B.116)
@@
Combining (B.104), (B.113), and (B.116) yields
$\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}\leq C\beta_{\mu}\left\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\right\|_{\mu^{\tau},\alpha^{\tau}}^{2},$
which concludes the proof. ∎